OpenStudy (anonymous):

A pound of nails cost $2. A pound of screws cost $6. How many pounds of nails are needed for a 5–pound mixture that costs $3.60 per pound?

OpenStudy (anonymous):

the options are 1,2,3,4

OpenStudy (ajprincess):

Let the number of pounds of nails be x and screws be y. x+y=5 x=5-y-(1) 2x+6y=5*3.6 2(5-y)+6y=18 10-2y+6y=18 10+4y=18 4y=18-10 4y=8 y=8/2 y=2 Nw substitute the value of y in (1) to find x.